# “It’s not a representation of me”: Examining Accent Bias and Digital Exclusion in Synthetic AI Voice Services

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 23–26,2025; Athens, Greececcs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Cultural characteristicsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In this study, we employ a mixed methods approach to study speech with diverse accents and pitch variations generated by two contemporary synthetic AI speech services (Speechify and ElevenLabs). We chose these services due to their significant market presence and distinct technical approaches to speech generation and voice cloning.
@@
The rest of the paper is organized as follows. In § 2, we present background and related work on accent representation and the current state of the art in synthetic AI voices. § 3 introduces the commercial platforms we evaluated in this study, followed by our methodology in § 4. We present quantitative results from our surveys in § 5 and qualitative insights from our interviews in § 6. Finally, in § 7, we conclude with actionable recommendations for stakeholders and discuss the limitations of our study .
@@ 2.1. Accent Representation and Belongingness in Media
Accents are a vital aspect of representation. However, academics and the media have found that diverse accents are underrepresented in the media and that this has negative emotional impacts on these speakers (Rangarajan, 2021; Mengesha et al., 2021). In fact, content analyses of popular media reveal how accents are often portrayed in ways that reflect and reinforce societal biases and stereotypes. Gluszek and Hansen (2016) found that accents were depicted as negative traits that hinder communication and underlie discrimination. Similarly, Levon et al. (2022) highlighted that “Non-standard” American and ”Foreign-Anglo” accents were associated with less favorable perceptions of status and physical traits in American primetime television. Dobrow and Gidney (1998) documented how accents were strategically used to differentiate between heroes and villains in children’s animated television programs. Even online polls ranking accent attractiveness reflect judgments that shape perceptions of speaker competence and social appeal (Bona, 2015; Jarvis, 2020; McLoone, 2024).
@@ 2.3. Synthetic AI Voices
Despite these advancements, synthetic AI voice services present several concerns. Hutiri et al. (2024) provide a comprehensive taxonomy of the harms associated with speech generators, emphasizing the urgent need for ethical guidelines and safeguards to mitigate these risks.
One challenge is the difficulty of distinguishing AI-generated voices from real voices. Studies show that TTS systems approach human-level quality compared to human recordings (Tan et al., 2024) and are nearly indistinguishable from human voices (Cambre et al., 2020). Synthetic AI voice clones also produce speech that may be mistaken for human voices (Barrington and Farid, 2024). Realistic, AI-generated voices raise concerns about impersonation, fraud, misinformation, privacy violations, and intellectual property theft (Bethea, 2024; Rose, 2024; Andrews, 2024).
@@ 3. Services
In this study, we primarily focus on the synthetic voice generation and voice cloning use cases. Of the various commercial providers offering these features, we investigate two synthetic AI voice services: Speechify and ElevenLabs. We chose them because they are popular—with 6 M and 23 M users, respectively, as of June 2024 (similarweb, 2024)—and they have a diverse selection of stock AI voices and voice cloning capabilities (see § A.1 for the services’ user interfaces).
Launched in 2017, Speechify was initially designed to help individuals access written content. Since then, the platform has expanded to other purposes such as reading for leisure and entertainment, as well as increasing productivity and efficiency for work-related tasks. Speechify has over 200 natural, life-like voices in over 60 languages and voice cloning capabilities in over 40 languages. Speechify claims to be the “#1 text to speech and AI voice generator” and they have “the most natural, human-sounding voice overs available on the market” (Speechify, 2024). Speechify notes they use an autoregressive voice model for their service.
ElevenLabs debuted in 2023 with the mission to “bridge language gaps, restore voices to those who have lost them, and make digital interactions feel more human, transforming the way we connect online”. They emphasize their realistic, human-like, and high-quality speech models in 32 languages, as well as voice cloning capabilities that purportedly capture human intonations and inflections (ElevenLabs, 2024a). ElevenLabs claims to focus on safe development, deployment, and usage of their services (ElevenLabs, [n. d.]). Synthetic AI voices from ElevenLabs were generated using their Eleven Multilingual v2 model (ElevenLabs, 2023), which was the default option at the time of our study.
@@ 4. Methodology
In this study, we conducted three experiments to evaluate the synthetic voice quality of stock AI voices and AI voice clones generated by Speechify and ElevenLabs. In addition, we interviewed native speakers to gather their perspectives on how their voices and accents were replicated by Speechify and ElevenLabs. In this section, we present the methods that we used in our three experiments. Figure 1 presents an overview of our study procedures. The mixed method design and procedure of our study were approved by the IRB of the authors’ university.
@@ 4.1.1. Text to Stock Synthetic AI Voice
The first goal of our study was to evaluate various aspects of the synthetic AI voices that are provided as stock options in these services. To achieve this, we generated 58 voices from English-language text with different accent and pitch combinations using the stock synthetic voice options from Speechify and ElevenLabs. We generated voices with high- and low-pitches. For each pitch, we generated voices with five accents: African, American, Australian, British, and Indian. We chose these parameters because (1) the pitches roughly correspond to masculine and feminine gendered voice registers, and (2) these five accents were available as options on ElevenLabs and Speechify. [^2] As of May 2024, ElevenLabs offered only those five accents while Speechify provided these same accents as options. When multiple voices were available for a given pitch and accent combination, we randomly selected three voices and generated voices from each one. For African accents, Speechify offered voices representing Kenya, Nigeria, and Tanzania, while ElevenLabs provided only Nigerian voices. Lastly, Speechify only offered one high-pitched Indian voice.
@@ 4.2.2. Survey Study Procedure
Similar to § 4.1, we recruited 303 crowdworkers from Prolific and followed the same inclusion criteria, which resulted in 250 crowdworkers. In this survey, however, participants listened to five pairs of recorded and cloned voices that matched the participants’ self-reported accent. One pair served as an attention check. As above, participants were not aware that they were listening to voices with the same accent as them. Further, they were also not informed which clip in each pair was the recorded or the cloned voice. For each participant, the order of voice pairs was randomized (i.e., Speechify or ElevenLabs first) and then the order of voices within each pair was randomized (i.e., cloned voice or recorded voice first). We anticipated that our task would take 10 minutes to complete and participants were compensated \$15 USD an hour.
@@ 5. Survey Results
In this section, we present our findings from Prolific crowdworkers’ evaluations of stock synthetic AI voices and AI voice clones generated by Speechify and ElevenLabs. Participant demographics can be found in § A.2. Results in § 5.1 and § 5.2 are from the approach we described in § 4.1, while the results in § 5.3 and § 5.4 are from the approach we described in § 4.2. We did not observe significant differences between high- and low-pitched voices across services, thus we present these results in § A.6.
@@ 5.1. Stock AI Voice Accent Prediction
We asked Prolific crowdworkers to predict the accent of the voice they were listening to without knowing they were listening to voices that matched their own accent, according to the services. 4(b) shows the mappings of the purported accents from each service to the predicted accents. 4(a) shows that participants agreed with the purported accents of Speechify’s American and British accents (98% and 90% of the time, respectively). However, there was less agreement with the purported accents of Speechify’s African, Australian, and Indian accents (78%, 81%, and 67% of the time, respectively). Participants listening to those accented voices interpreted them as having British or American accents.
4(b) shows that participants disagreed with ElevenLabs accent designations more often than Speechify. Like with Speechify, participants tended to agree with the purported accents of ElevenLabs’ American and British accents (100% and 82%, respectively), but the vast majority of participants disagreed with the designations of Indian and African accented voices (only 12% and 10% agreed, respectively). All participants disagreed with the designations of Australian accented voices.
@@ 5.2. Stock AI Voice Quality
We aggregated the MOS by accent and service as seen in Figure 6. A score of 4.3–4.5 indicates excellent quality (Words, 2023; Twilio, 2024) and we see that no accents across the two services meet this baseline. While Speechify produced voices with relatively consistent MOS across all accents, their ratings were less than those for ElevenLabs. However, it is important to qualify ElevenLabs’ higher ratings for African, Indian, and Australian accents with the fact that almost all of our participants thought they were listening to American or British accents. For an analysis of stock AI voice quality when participants’ accent labels aligned with the services’ labels, see § A.5.
We used a linear mixed model to quantify the main effects of accent, service, and pitch on voice quality, while adjusting for random rater identity and audio clip effects. Figure 6 shows that both Speechify voices and low-pitched voices have significantly lower MOS (-0.29 and -0.28 points, respectively). We do not observe significant differences across accents.
@@ 5.3. Naturalness of AI Voice Clones Across Services
Across both services, we observe no clear trend that one voice was perceived as having greater naturalness versus the other. 7(b)a shows that AI-generated cloned voices with African and British accents from Speechify were deemed as more natural by participants. The ElevenLabs AI cloned voices with American and British accents were judged as sounding more natural than the recorded voices, as shown in 7(b)b. We ran a 1-way ANOVA to check whether there is a significant variation in the proportion of participants that preferred the voice clone across accents and we did not find any significant variation from either service.
@@ 5.4. Accent Mimicry Accuracy of AI Voice Clones Across Services
Figure 8 illustrates the AMA of the AI cloned voices from each service with respect to the original voices. Both services had relatively consistent AMA across all accents, however, the AI cloned voices from ElevenLabs were judged as having slightly greater accuracy overall than the cloned voices from Speechify. The average AMA for Speechify is 69% and for ElevenLabs is 78.5%. These averages were significantly different from a t-test that resulted in a $t=6.6422$ and $p<0.05$.
@@ 6.1.3. Understandability and Interaction Challenges
(b) Voice Clone Accent Mimicry Quality
@@ 6.2.3. Participant’s Practical Use Cases and Concerns about the Landscape
Overall, participants expressed a range of concerns which included issues related to privacy, scams, potential misuse, lack of regulation and safeguards, misrepresentation, and broader ethical implications. These concerns align with Hutiri et al’s taxonomy of consequential harms from speech generators (Hutiri et al., 2024). Some participants also voiced political concerns, particularly because the majority of interviews were conducted prior to the 2024 U.S. presidential election which heightened anxieties about potential misuses of AI voices. Interestingly, ElevenLabs’ more realistic voice clones seemed to amplify these concerns while the clones generated by Speechify generally caused fewer or no concerns. P2 concludes, “I think, yeah, my takeaway would be be afraid. Be very afraid, but not of the second one [from Speechify]”. However, not all participants shared these concerns. P17 expressed confidence that worries surrounding AI speech technologies were typical of any emerging technology and would likely diminish with time, analogizing “I’m accepting, that’s the price we pay…, People who chose to stay with horse and cart were put at a disadvantage.
@@ 7. Discussion
In this study, we conducted an in-depth examination of accent bias and digital exclusion in synthetic AI voice services through a mixed-methods analysis of Speechify and ElevenLabs. Our investigation of these two prominent services reveals insights into the current state of AI voice technology. While Speechify demonstrated less accent dilution (a.k.a. accent leveling) for non-American and non-British speakers, ElevenLabs achieved greater accuracy in capturing accent characteristics during voice cloning. However, our research shows that this accent-specific technical performance is only one dimension of how users experience and evaluate these technologies.
Our analysis presents three perspectives on synthetic AI voice services. First, our quantitative evaluations reveal measurable differences in how these systems perform across various accents and pitch combinations. Second, our user preference data demonstrates that accent accuracy alone does not determine user satisfaction with voice clones, suggesting a more complex relationship between technical performance and user acceptance. Third, our qualitative findings uncover the deeper emotional and cultural dimensions of how users interact with AI-generated versions of their own voices.
@@ 7.2. Limitations and Future Work
Our study evaluated Speechify and ElevenLabs as representative examples of current commercial AI voice technology. Although our study captured relevant user sentiments and technical capabilities using these two services, future work can extend our study by examining a broader range of synthetic voice systems. We used the modified MOS measure (Viswanathan and Viswanathan, 2005), which does not have a validated baseline, and therefore we used the industry MOS measure’s baseline. Future work could develop and validate a baseline for this modified MOS. In our quantitative evaluation of AI voice clones, we focused on accent accuracy and naturalness metrics, but did not explicitly assess participant preferences between original and cloned voices or their ability to identify AI-generated speech. This limitation may have constrained our understanding of how users perceive authenticity in synthetic voices.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732018isbn: 979-8-4007-1482-5/2025/06ccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Cultural characteristicsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In this study, we employ a mixed methods approach to study speech with diverse English-language accents and pitch variations generated by two contemporary synthetic AI speech services (Speechify and ElevenLabs). We chose these services due to their significant market presence and distinct technical approaches to speech generation and voice cloning.
@@ 2.1. Accent Representation and Belongingness in Media
Accents are a vital aspect of representation. In linguistics, accent refers to patterns of pronunciation, distinguishing it from dialect, which also includes grammar and vocabulary. While everyone has an accent, everyday usage often treats certain ways of speaking as “accented” and others as neutral. Accents are shaped by geography, class, ethnicity, and culture, and perceptions of them are socially and ideologically loaded (Markl and Lai, 2023).
However, academics and the media have found that diverse accents are underrepresented in the media and that this has negative emotional impacts on these speakers (Rangarajan, 2021; Mengesha et al., 2021). In fact, content analyses of popular media reveal how accents are often portrayed in ways that reflect and reinforce societal biases and stereotypes. Gluszek and Hansen (2016) found that accents were depicted as negative traits that hinder communication and underlie discrimination. Similarly, Levon et al. (2022) highlighted that “Non-standard” American and ”Foreign-Anglo” accents were associated with less favorable perceptions of status and physical traits in American primetime television. Dobrow and Gidney (1998) documented how accents were strategically used to differentiate between heroes and villains in children’s animated television programs. Even online polls ranking accent attractiveness reflect judgments that shape perceptions of speaker competence and social appeal (Bona, 2015; Jarvis, 2020; McLoone, 2024).
@@ 2.3. Synthetic AI Voices
Despite these advancements, synthetic AI voice services present several concerns. Székely et al. (Székely et al., 2025) argue that the socioindexical properties of AI-generated voices can influence users’ speech patterns and identity expression, potentially reinforcing normative biases by defaulting to voices that are most commonly used and culturally dominant. Hutiri et al. (2024) provide a comprehensive taxonomy of the harms associated with speech generators, emphasizing the urgent need for ethical guidelines and safeguards to mitigate these risks.
One challenge is the difficulty of distinguishing AI-generated voices from real voices. Studies show that TTS systems approach human-level quality compared to human recordings (Tan et al., 2024) and are nearly indistinguishable from human voices (Cambre et al., 2020). Synthetic AI voice clones also produce speech that may be mistaken for human voices (Barrington et al., 2025). Realistic, AI-generated voices raise concerns about impersonation, fraud, misinformation, privacy violations, and intellectual property theft (Bethea, 2024; Rose, 2024; Andrews, 2024).
Another challenge is the lack of specificity in how AI-generated voice systems label accents, particularly varieties of English. These accent labels are often ambiguous. The way voice service providers describe accents often does not align with linguistic definitions, and instead may reflect broader or regional classifications  (Roberts, 2021; Bender, 2019).
@@ 3. Services
In this study, we primarily focus on the synthetic voice generation and voice cloning use cases in the English language. Of the various commercial providers offering these features, we investigate two synthetic AI voice services: Speechify and ElevenLabs. We chose them because they are popular—with 6 M and 23 M users, respectively, as of June 2024 (similarweb, 2024)—and they have a diverse selection of stock AI voices and voice cloning capabilities (see § A.1 for the services’ user interfaces).
Launched in 2017, Speechify was initially designed to help individuals access written content. Since then, the platform has expanded to other purposes such as reading for leisure and entertainment, as well as increasing productivity and efficiency for work-related tasks. Speechify has over 200 “natural, life-like” voices in over 60 languages and voice cloning capabilities in over 40 languages. Speechify claims to be the “#1 text to speech and AI voice generator” and they have “the most natural, human-sounding voice overs available on the market” (Speechify, 2024). Speechify notes they use an autoregressive voice model for their service.
ElevenLabs debuted in 2023 with the mission to “bridge language gaps, restore voices to those who have lost them, and make digital interactions feel more human, transforming the way we connect online”. They emphasize their ‘realistic, human-like, and high-quality” speech models in 32 languages, as well as voice cloning capabilities that purportedly capture “human intonations and inflections” (ElevenLabs, 2024a). ElevenLabs claims to focus on safe development, deployment, and usage of their services  (ElevenLabs, [n. d.]). Synthetic AI voices from ElevenLabs were generated using their Eleven Multilingual v2 model  (ElevenLabs, 2023), which was the default option at the time of our study.
@@ 4. Methodology
In this study, we conducted three experiments to evaluate the synthetic voice quality of English-language stock AI voices and AI voice clones [^2] https://huggingface.co/datasets/shiramichel/ai-speech-biases generated by Speechify and ElevenLabs. In addition, we interviewed native speakers to gather their perspectives on how their voices and accents were replicated by Speechify and ElevenLabs. In this section, we present the methods that we used in our three experiments. Figure 1 presents an overview of our study procedures. The mixed method design and procedure of our study were approved by the IRB of the authors’ university.
@@ 4.1.1. Text to Stock Synthetic AI Voice
The first goal of our study was to evaluate various aspects of the synthetic AI voices that are provided as stock options in these services. To achieve this, we generated 58 voices from English-language text with different accent and pitch combinations using the stock synthetic voice options from Speechify and ElevenLabs. We generated voices with high- and low-pitches. For each pitch, we generated voices with five accents: African, American, Australian, British, and Indian. We chose these parameters because (1) the pitches roughly correspond to masculine and feminine gendered voice registers, and (2) these five accents were available as options on ElevenLabs and Speechify. [^3] As of May 2024, ElevenLabs offered only those five accents while Speechify provided these same accents as options. It is also important to note that these accent labels can be ambiguous, often encompassing a wide range of distinct regional variations. When multiple voices were available for a given pitch and accent combination, we randomly selected three voices and generated voices from each one. For African accents, Speechify offered voices representing Kenya, Nigeria, and Tanzania, while ElevenLabs provided only Nigerian voices. Lastly, Speechify only offered one high-pitched Indian voice.
@@ 4.2.2. Survey Study Procedure
Similar to § 4.1, we recruited 303 crowdworkers from Prolific and followed the same inclusion criteria, which resulted in 250 crowdworkers. In this survey, however, participants listened to five pairs of recorded and cloned voices that matched the participants’ self-reported accent. One pair served as an attention check. As above, participants were not aware that they were only listening to voices with the same accent as them. Further, they were also not informed which clip in each pair was the recorded or the cloned voice. For each participant, the order of voice pairs was randomized (i.e., Speechify or ElevenLabs first) and then the order of voices within each pair was randomized (i.e., cloned voice or recorded voice first). We anticipated that our task would take 10 minutes to complete and participants were compensated \$15 USD an hour.
@@ 5. Survey Results
In this section, we present our findings from Prolific crowdworkers’ evaluations of English-language stock synthetic AI voices and AI voice clones generated by Speechify and ElevenLabs. Participant demographics can be found in § A.2. Results in § 5.1 and § 5.2 are from the approach we described in § 4.1, while the results in § 5.3 and § 5.4 are from the approach we described in § 4.2. We did not observe significant differences between high- and low-pitched voices across services, thus we present these results in § A.7.
@@ 5.1. Stock AI Voice Accent Prediction
We asked Prolific crowdworkers to predict the English-language accent of the voice they were listening to without knowing they were listening to voices that matched their own accent, according to the services. 4(b) shows the mappings of the purported accents from each service to the predicted accents. 4(a) shows that participants agreed with the purported accents of Speechify’s American and British accents (98% and 90% of the time, respectively). However, there was less agreement with the purported accents of Speechify’s African, Australian, and Indian accents (78%, 81%, and 67% of the time, respectively). Participants listening to those accented voices interpreted them as having British or American accents.
4(b) shows that participants disagreed with ElevenLabs accent designations more often than Speechify. Like with Speechify, participants tended to agree with the purported accents of ElevenLabs’ American and British accents (100% and 82%, respectively), but the vast majority of participants disagreed with the designations of Indian and African accented voices (only 12% and 10% agreed, respectively). All participants disagreed with the designations of Australian accented voices. These findings suggest that while AI-generated American and British accents are largely perceived as intended, labels for African, Indian, and Australian voices often mismatch listener perceptions, highlighting the need for more accurate and culturally informed accent designations.
@@ 5.2. Stock AI Voice Quality
We aggregated the MOS by accent and service as seen in Figure 5. A score of 4.3–4.5 indicates excellent quality (Words, 2023; Twilio, 2024) and we see that no English-language accents across the two services meet this baseline. While Speechify produced voices with relatively consistent MOS across all accents, their ratings were less than those for ElevenLabs. However, it is important to qualify ElevenLabs’ higher ratings for African, Indian, and Australian accents with the fact that almost all of our participants thought they were listening to American or British accents. For an analysis of stock AI voice quality when participants’ accent labels aligned with the services’ labels, see § A.5.
We used a linear mixed model to quantify the main effects of accent, service, and pitch on voice quality, while adjusting for random rater identity and audio clip effects. Figure 6 shows that both Speechify voices and low-pitched voices have significantly lower MOS (-0.29 and -0.28 points, respectively). We do not observe significant differences across accents. For an analysis of the linear mixed model with respect to each service, see § A.6.
Although ElevenLabs voices received higher quality ratings than Speechify, these ratings may have been influenced by mismatches between how participants perceived the accents and how the services labeled them. This suggests that listener perceptions have a significant impact on how the quality of synthetic voices is evaluated.
@@ 5.3. Naturalness of AI Voice Clones Across Services
Across both services, we observe no clear trend that one voice was perceived as having greater naturalness versus the other. 7(b)a shows that AI-generated cloned voices with African and British accents from Speechify were deemed as more natural by participants. The ElevenLabs AI cloned voices with American and British accents were judged as sounding more natural than the recorded voices, as shown in 7(b)b. We ran a 1-way ANOVA to check whether there is a significant variation in the proportion of participants that preferred the voice clone across accents and we did not find any significant variation from either service. These results suggest that perceptions of naturalness in AI-cloned voices vary by accent but do not consistently favor either recorded or synthetic speech, indicating there is a more intricate connection between technical performance and user acceptance.
@@ 5.4. Accent Mimicry Accuracy of AI Voice Clones Across Services
Figure 8 illustrates the AMA of the AI cloned voices from each service with respect to the original voices. Both services had relatively consistent AMA across all accents, however, the AI cloned voices from ElevenLabs were judged as having slightly greater accuracy overall than the cloned voices from Speechify. The average AMA for Speechify is 69% and for ElevenLabs is 78.5%. These averages were significantly different from a t-test that resulted in a $t=6.6422$ and $p<0.05$. The findings suggest that while both services show consistent accent mimicry, ElevenLabs voices are more often perceived as accurately capturing the original English-language accent, reflecting stronger performance in accent replication.
@@ 6.1.3. Understandability and Interaction Challenges
(b) Voice Clone Accent Mimicry Quality (AMQ)
@@ 6.2.3. Participant’s Practical Use Cases and Concerns about the Landscape
Overall, participants expressed a range of concerns which included issues related to privacy, scams, potential misuse, lack of regulation and safeguards, misrepresentation, and broader ethical implications. These concerns align with Hutiri et al’s taxonomy of consequential harms from speech generators (Hutiri et al., 2024). Some participants also voiced political concerns, particularly because the majority of interviews were conducted prior to the 2024 U.S. presidential election which heightened anxieties about potential misuses of synthetic AI voices. Interestingly, ElevenLabs’ more realistic voice clones seemed to amplify these concerns while the clones generated by Speechify generally caused fewer or no concerns. P2 concludes, “I think, yeah, my takeaway would be be afraid. Be very afraid, but not of the second one [from Speechify]”. However, not all participants shared these concerns. P17 expressed confidence that worries surrounding AI speech technologies were typical of any emerging technology and would likely diminish with time, analogizing “I’m accepting, that’s the price we pay…, People who chose to stay with horse and cart were put at a disadvantage.
@@ 7. Discussion
In this study, we conducted an in-depth examination of English-langauge accent bias and digital exclusion in synthetic AI voice services through a mixed-methods analysis of Speechify and ElevenLabs. Our investigation of these two prominent services reveals insights into the current state of AI voice technology. While Speechify demonstrated less accent dilution (a.k.a. accent leveling) for non-American and non-British speakers, ElevenLabs achieved greater accuracy in capturing accent characteristics during voice cloning. However, our research shows that this accent-specific technical performance is only one dimension of how users experience and evaluate these technologies.
Our analysis presents three perspectives on synthetic AI voice services. First, our quantitative analysis shows that while crowdworker participants often disagreed with the classification of voices with Indian, Australian, and African accents, especially those generated by ElevenLabs, both stock AI voices and AI-cloned voices from ElevenLabs were evaluated as having higher voice and accent quality. Interview participants also generally perceived the ElevenLabs’ clones as more successful in recreating their voice. Second, our user preference data demonstrates that accent accuracy alone does not determine user satisfaction with voice clones. Naturalness ratings by crowdworkers, combined with interview participants’ lack of enthusiasm for practical use cases and concerns about ethical risks, suggest a more complex relationship between technical performance and user acceptance. In particular, the high realism of ElevenLabs’ clones amplified anxieties for many. Third, our qualitative findings uncover the deeper emotional and cultural dimensions of how users interact with AI-generated versions of their own voices. Participants noted that most AI systems privilege American and British accents, and those without these accents often shared feelings of exclusion and misrepresentation.
@@ 7.2. Limitations and Future Work
Our study evaluated Speechify and ElevenLabs as representative examples of current commercial AI voice technology. Although our study captured relevant user sentiments and technical capabilities using these two services, future work can extend our study by examining a broader range of synthetic voice systems. A major limitation of our study is that its findings are limited to English, and may not generalize to lower-resource or non-English languages. Another limitation of our study is that the Speech Accent Archive  (Weinberger and Kunath, 2011) primarily consists of individuals who are immigrants, and as a result, birth country may not be a reliable descriptor of accent. We recognize that immigration can influence one’s accent in ways that birth country alone does not capture. While we did not initially consider immigration status as a potential confounding factor, our decision to filter by birth country aligns with our approach to filtering Prolific crowdworkers based on “place of most time spent before turning 18” to ensure that participants listened to accented voices that are likely to be familiar with their own accent. We also used the modified MOS measure (Viswanathan and Viswanathan, 2005), which does not have a validated baseline, and therefore we used the industry MOS measure’s baseline. Future work could develop and validate a baseline for this modified MOS. In our quantitative evaluation of AI voice clones, we focused on accent accuracy and naturalness metrics, but did not explicitly assess participant preferences between original and cloned voices or their ability to identify AI-generated speech. This limitation may have constrained our understanding of how users perceive authenticity in synthetic voices.